OpenStudy (anonymous):

8√y + 9√y add simplify

OpenStudy (unklerhaukus):

\[8\sqrt y+9 \sqrt y=(8+9)\sqrt y\]

OpenStudy (unklerhaukus):

\[=17\sqrt y\]